OpenStudy (anonymous):

A bottle contains 400 mL of fruit punch with a concentration of 50% pure fruit juice. Jill drinks 100 mL of the punch and then refills the bottle with an equal amount of a cheaper brand of punch. If the concentration of juice in the bottle is now reduced to 48%, what was the concentration in the punch that Jill added?

OpenStudy (anonymous):

make a chart

OpenStudy (anonymous):

that doesnt help me do you know another way to solve it

OpenStudy (anonymous):

it does help

OpenStudy (anonymous):

o it doesnt help you, ok let me try another way

OpenStudy (anonymous):

Jill drinks 100 ml of the punch: Now, the volume is 400-100 = 300 ml The amount of fruit juice is now: .50*300 = 150 ml

OpenStudy (anonymous):

so what is the percentage of pure fruit juice

OpenStudy (anonymous):

100x+150=0.48*(400) 100x+150=192 x= (192-150)/100 x=0.42

OpenStudy (anonymous):

if you like my answer click good answer,ty